OpenStudy (anonymous):

If f(x) varies directly with x^2, and f(x) = 72 when x = 6, find the value of f(3). A. 108 B. 18 C. 36 D. 54

OpenStudy (anonymous):

f(x)=kx^2 f(6)=k36=72 so k=2 then : f(x)=2x^2 and f(3)=2*9=18